I am 30 years old and my sister is 10. I need an equation to figure out how old my sister and I will be when I am three times as old as my sister.
Answer by stanbon(75887)

I am 30 years old and my sister is 10. I need an equation to figure out how old my sister and I will be when I am three times as old as my sister.
----
You are now 3 times as old as old as your sister.
That will never happen at a later or earlier age.
-----
Maybe you would like to know whan you will be 5 times
as old as your sister.
Equation:
(30+x) = 5(10+x)
30+x = 50 + 5x
4x = -20
x = -5
----
When you are 30-5 = 25 you will be 5 times as old
as your sister, who will be 10-5 = 5...
So that happened in the past; and it will never happen again.
Cheers,
Stan H.